OpenStudy (anonymous):

45 = 3b + 69 sole the equation show your work ???

OpenStudy (anonymous):

solve***

OpenStudy (anonymous):

45=3b+69 45-69=3b -24=3b b=-24/3 b=8

OpenStudy (anonymous):

y am i so confused! 8x – 12 = 4x + 24

OpenStudy (anonymous):

you need to move all the x's to one side of the equation

OpenStudy (anonymous):

8x-4x=24+12 4x=36 x=9